OpenStudy (anonymous):

What is the height of the triangle if the area is sixty square meters? A triangle with a base labeled twelve meters and a second side labeled fourteen meters. A dotted line extends from the vertex downward and perpendicular to the base. This dotted line is not labeled. 9 meters ten meters nineteen meters twenty meters

OpenStudy (anonymous):

Please help i will fan and medal

ganeshie8 (ganeshie8):

use Area of triangle formula

ganeshie8 (ganeshie8):

\(A = \frac{1}{2} bh\)

ganeshie8 (ganeshie8):

you're given \(A = 60\), \(b = 12\) plug them in the formula and solve \(h\)

ganeshie8 (ganeshie8):

\(60 = \frac{1}{2}*12*h\)

ganeshie8 (ganeshie8):

solve \(h\)

OpenStudy (anonymous):

12

ganeshie8 (ganeshie8):

no, try again

OpenStudy (anonymous):

What is H

ganeshie8 (ganeshie8):

\(60 = \frac{1}{2}*12*h\) \(60 = 6h\) \(? = h\)

ganeshie8 (ganeshie8):

the link wont open for me (non flvs ppl)

ganeshie8 (ganeshie8):

did u solve \(h\) yet ? :)